whats the area of a rectangle that has the length of 3" and a width of 2" ?
Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

A=LW
A=3*2=6^2" ANS.

Answer by Alan3354(69443)

2 inches * 3 inches = 6 square inches